OpenStudy (anonymous):

Y=x^2-49 factor the expression

OpenStudy (henryblah):

Do you remember difference of 2 squares?

OpenStudy (jdoe0001):

\(\bf y=x^2-49 \qquad \qquad 49=7^2\qquad thus\\ \quad \\ y = x^2-7^2\\ \quad \\ \textit{recall that }\qquad a^2-b^2 = (a-b)(a+b)\)